OpenStudy (anonymous):

solve the following equation for x 23x-17y=5

imqwerty (imqwerty):

23x=5+17y x=(5+17y)/23

OpenStudy (anonymous):

add 17y to both sides

OpenStudy (anonymous):

im not sure if i got this right I'm really bad at math but I'm trying wouldn't it cancel out on the left side when i add 17y and when i add 17y to the right side i got 40 ?? @imqwerty @surjithayer